# GMAT: Grounded Multi-Agent Clinical Description Generation for Text Encoder in Vision-Language MIL for Whole Slide Image Classification

### v1

@@ Abstract
Keywords:
Multi-Agent Systems Whole Slide Images (WSIs) Multiple Instance Learning (MIL) Vision-Language Models (VLM)
@@ 1 Introduction
To address the scale and complexity of WSIs, Multiple Instance Learning (MIL) has become the dominant approach for weakly supervised classification. In this framework, a slide is treated as a bag of image patches (instances), with supervision provided only at the slide level. Early models like ABMIL  [1], CLAM  [2], and TransMIL  [3] leverage attention- or transformer-based aggregation to summarize patch features effectively. More recent work introduces advanced architectures, including HIPT  [4] with hierarchical transformers, DSMIL  [5] with dual-stream learning, and others such as CAMIL [6], DTFD-MIL  [7], SNUFFY [8], and DGMIL [9], each proposing novel strategies for aggregation or feature modeling to improve classification performance.
Vision-Language Models (VLMs) such as CLIP have been incorporated into MIL pipelines to improve classification and interpretability by aligning visual features with textual prompts like disease labels or descriptive phrases. A common strategy involves prompt tuning, using either handcrafted or LLM-generated prompts. MGPath  [10] introduces a multi-granular prompt framework that adapts across tissue magnifications to support few-shot learning. ViLaMIL  [11] integrates magnification-aware embeddings with hierarchical attention across patch scales. MSCPT  [12] proposes a multi-scale, context-aware prompt tuning method that aligns vision-text embeddings across magnifications. These approaches highlight the potential of multiscale integration and prompt design in enhancing VLM-based WSI classification. These works underscore the role of prompt design and scale integration in advancing VLM-based digital pathology. Building on this, recent efforts are now leveraging foundation models to further extend vision-language and vision-only capabilities.
Recent progress in Whole Slide Imaging (WSI) has been driven by Vision-Language Models (VLMs) like BioCLIP [13], PLIP [14], and CONCH [15], which align images with text for zero-shot performance. Dual-scale models such as ViLa-MIL [16] enhance fine-grained reasoning. In parallel, vision-only models like CTransPath [17] and GigaPath [18] learn strong representations without text, enabling scalable classification in low-label settings.
@@ 3.2 Methods Under Comparision
We evaluate zero-shot and fine-tuned performance across multiple settings. For zero-shot, we assess CONCH using either a single-description setup (denoted as Single Class Description in the table) or a list of GMATG-generated descriptions. The single-description setup follows the two-level prompt format from ViLa-MIL [16], covering 5× and 10× magnifications. For fine-tuning, we compare ViLa-MIL with standard prompts to our GMAT framework. All methods use the same CONCH encoder for both text and image to ensure fairness. An ablation study further examines the impact of the multi-agent design versus a single-agent variant.
@@ 3.3 Result and Analysis
First, we evaluate the effectiveness of our approach in a zero-shot setting using the CONCH model. Specifically, we compare performance using a single class description versus a list of descriptions generated by GMATG. For the class description baseline, we adopt the two-level prompt structure used in ViLa-MIL  [16]. In contrast, the list-based variant includes multiple, structured descriptions per class, covering diverse pathological aspects. Results show that the list-based approach consistently improves AUC, F1 score across both TCGA-RCC and TCGA-Lung datasets.
@@
CONCH Single Class Description 0.5730 $\pm$± 0.0314 0.3466 $\pm$± 0.0138 0.4821 $\pm$± 0.0070
CONCH List from GMATG 0.5912 $\pm$± 0.0328 0.3691 $\pm$± 0.0347 0.4357 $\pm$± 0.0514
@@
ViLa-MIL Single Class Description 0.9844 $\pm$± 0.0070 0.9028 $\pm$± 0.0445 0.9197 $\pm$± 0.0184
GMAT List from GMATG 0.9791 $\pm$± 0.0116 0.9131 $\pm$± 0.0293 0.9262 $\pm$± 0.0294
@@
CONCH Single Class Description 0.6767 $\pm$± 0.0288 0.6116 $\pm$± 0.0164 0.6300 $\pm$± 0.0089
CONCH List from GMATG 0.7226 $\pm$± 0.0233 0.6693 $\pm$± 0.0262 0.6711 $\pm$± 0.0247
@@
ViLa-MIL Single Class Description 0.9499 $\pm$± 0.0308 0.8894 $\pm$± 0.0418 0.8899 $\pm$± 0.0422
GMAT List from GMATG 0.9641 $\pm$± 0.0057 0.9023 $\pm$± 0.0184 0.9028 $\pm$± 0.0183

### v2

@@ 1 Introduction
To address the scale and complexity of WSIs, Multiple Instance Learning (MIL) has become the dominant approach for weakly supervised classification. In this framework, a slide is treated as a bag of image patches (instances), with supervision provided only at the slide level. Early models like ABMIL  [ilse2018attention-abmil], CLAM  [lu2021data-clam], and TransMIL  [shao2021transmil] leverage attention- or transformer-based aggregation to summarize patch features effectively. More recent work introduces advanced architectures, including HIPT  [chen2022scaling-hipt] with hierarchical transformers, DSMIL  [li2021dual-dsmil] with dual-stream learning, and others such as CAMIL [camil_fourkioti2024camil], DTFD-MIL  [dtfd_zhang_dtfd-mil_2022], SNUFFY [snuffy_afarinia2024snuffyefficientslideimage], and DGMIL [qu2022dgmil], each proposing novel strategies for aggregation or feature modeling to improve classification performance.
Vision-Language Models (VLMs) such as CLIP have been incorporated into MIL pipelines to improve classification and interpretability by aligning visual features with textual prompts like disease labels or descriptive phrases. A common strategy involves prompt tuning, using either handcrafted or LLM-generated prompts. MGPath  [nguyen2025mgpathvisionlanguagemodelmultigranular] introduces a multi-granular prompt framework that adapts across tissue magnifications to support few-shot learning. ViLaMIL  [nguyen2025fewshot] integrates magnification-aware embeddings with hierarchical attention across patch scales. MSCPT  [10979677] proposes a multi-scale, context-aware prompt tuning method that aligns vision-text embeddings across magnifications. These approaches highlight the potential of multiscale integration and prompt design in enhancing VLM-based WSI classification. These works underscore the role of prompt design and scale integration in advancing VLM-based digital pathology. Building on this, recent efforts are now leveraging foundation models to further extend vision-language and vision-only capabilities.
Recent progress in Whole Slide Imaging (WSI) has been driven by Vision-Language Models (VLMs) like BioCLIP [zhang2024biomedclip], PLIP [huang2023plip], and CONCH [lu2023mizero], which align images with text for zero-shot performance. Dual-scale models such as ViLa-MIL [shi2024vilamil] enhance fine-grained reasoning. In parallel, vision-only models like CTransPath [wang2022ssltransformer] and GigaPath [xu2024slidefound] learn strong representations without text, enabling scalable classification in low-label settings.
@@ 3.2 Methods Under Comparision
We evaluate zero-shot and fine-tuned performance across multiple settings. For zero-shot, we assess CONCH using either a single-description setup (denoted as Single Class Description in the table) or a list of GMATG-generated descriptions. The single-description setup follows the two-level prompt format from ViLa-MIL [shi2024vilamil], covering 5× and 10× magnifications. For fine-tuning, we compare ViLa-MIL with standard prompts to our GMAT framework. All methods use the same CONCH encoder for both text and image to ensure fairness. An ablation study further examines the impact of the multi-agent design versus a single-agent variant.
@@ 3.3 Result and Analysis
First, we evaluate the effectiveness of our approach in a zero-shot setting using the CONCH model. Specifically, we compare performance using a single class description versus a list of descriptions generated by GMATG. For the class description baseline, we adopt the two-level prompt structure used in ViLa-MIL  [shi2024vilamil]. In contrast, the list-based variant includes multiple, structured descriptions per class, covering diverse pathological aspects. Results show that the list-based approach consistently improves AUC, F1 score across both TCGA-RCC and TCGA-Lung datasets.
@@
CONCH Single Class Description 0.5730 $\pm$ 0.0314 0.3466 $\pm$ 0.0138 0.4821 $\pm$ 0.0070
CONCH List from GMATG 0.5912 $\pm$ 0.0328 0.3691 $\pm$ 0.0347 0.4357 $\pm$ 0.0514
@@
ViLa-MIL Single Class Description 0.9844 $\pm$ 0.0070 0.9028 $\pm$ 0.0445 0.9197 $\pm$ 0.0184
GMAT List from GMATG 0.9791 $\pm$ 0.0116 0.9131 $\pm$ 0.0293 0.9262 $\pm$ 0.0294
@@
CONCH Single Class Description 0.6767 $\pm$ 0.0288 0.6116 $\pm$ 0.0164 0.6300 $\pm$ 0.0089
CONCH List from GMATG 0.7226 $\pm$ 0.0233 0.6693 $\pm$ 0.0262 0.6711 $\pm$ 0.0247
@@
ViLa-MIL Single Class Description 0.9499 $\pm$ 0.0308 0.8894 $\pm$ 0.0418 0.8899 $\pm$ 0.0422
GMAT List from GMATG 0.9641 $\pm$ 0.0057 0.9023 $\pm$ 0.0184 0.9028 $\pm$ 0.0183